Find the acceleration of a car a velocity of 60 m/s in a distance 360 m

1
Expert's answer
2022-03-02T10:06:50-0500

Explanations & Calculations


  • Assuming it started at rest and travelled 360m before gaining that velocity 60m/s at constant acceleration, apply "v^2=u^2+2as" to calculate the acceleration.

"\qquad\qquad
\begin{aligned}
\small (60\,ms^{-1})^2&=\small 0^2+2a(360\,m)\\
\small a&=\small 5\,ms^{-2}
\end{aligned}"